OpenStudy (anonymous):

Pablo rolled a standard die 48 times. He got a 1 two times. How does this result compare to the expected results? Explain.

OpenStudy (anonymous):

you "expect" one sixth of the rolls to be a 1, because the probability you get a 1 is 1/6 \[\frac{1}{6}\times 48=8\] so you would expect 8 ones, not 2

OpenStudy (anonymous):

I would explain that Pablo got very lucky (or unlucky)